# PyramidKV: Dynamic KV Cache Compression based on Pyramidal Information Funneling

### v2

@@ 5.3 Baselines
StreamingLLM (SLM)
[32] is an efficient framework that enables LLMs trained with a finite length attention window to generalize to infinite sequence length without any fine-tuning. They propose StreamingLLM based on the attention sink phenomenon that keeping the KV of the first few tokens will largely recover the performance of window attention. StreamingLLM is a competitive method to solve long-context tasks. In our experiments, to be consistent with other methods, we simply keep the KV cache of the last $\alpha$ tokens and the first $k-\alpha$ tokens, as suggested in the paper.
Heavy Hitter Oracle (H2O)
[34] is a KV cache compression policy that dynamically retains a balance of recent and Heavy Hitter (H2) tokens. H2O keeps a fixed cache size of the Key and Value matrix across Transformer layers. The selection process for the KV cache is driven by attention scores, specifically utilizing the average attention scores from all queries across all tokens to guide the selection.
@@
[22] automatically compresses KV caches by selecting clustered important tokens for each attention head. This method discerns the attention patterns of the Key and Value matrices using a localized observation window positioned at the end of the prompts. However, unlike H2O, SnapKV employs a more nuanced clustering algorithm that includes a pooling layer. Additionally, SnapKV captures attention signals using patterns from a localized window (Instruction Tokens), rather than aggregating attention across all queries, allowing for more targeted and efficient compression.
@@
5.4 Experiment Setup
We maintain a fixed constant KV cache size for each layer for the baseline methods. In contrast, PyramidKV employs varying KV cache sizes across different layers. To ensure a fair comparison, we adjusted the average KV cache size in PyramidKV to match that of the baseline models, to keep the total memory consumption of all methods the same. In our experiment, we set $\beta=20$ and $\alpha=8$. We use the same prompt for each dataset in all the experiments.
5.5 PyramidKV Excels in Memory-constrained Settings
Figure 4: The evaluation results from LongBench[2] across 64, 96, 128, 256 and 512 cache sizes. The evaluation metrics are the average score of LongBench across datasets.  PyramidKV outperforms H2O [34], SnapKV [22] and StreamingLLM [32], especially in small KV cache sizes.
@@ 5.5 PyramidKV Excels in Memory-constrained Settings
LlaMa-3-8B-Instruct, KV Size = 128
@@
FKV 25.43 31.72 48.31 42.21 26.78 17.50 25.45 23.90 4.98 68.50 86.33 42.44 5.00 88.40 51.30 47.89 39.76
Mistral-7B-Instruct, KV Size = 128
SKV 19.16 21.46 43.52 38.60 23.35 16.09 17.66 21.84 3.89 47.50 84.15 40.24 5.00 69.31 46.98 42.97 33.85
H2O 21.20 21.90 41.55 33.56 21.28 12.93 18.59 22.61 3.97 39.00 82.37 40.44 6.00 83.19 46.41 42.66 33.60
SLM 16.57 14.68 32.40 30.19 22.64 12.34 18.08 18.96 3.71 43.50 74.22 29.02 4.50 29.48 39.23 36.16 26.60
Ours 21.75 22.03 44.32 34.06 22.79 15.77 18.58 21.89 4.03 66.00 83.46 39.75 4.50 66.90 46.96 41.28 34.62
@@
SKV 25.89 32.93 48.56 42.96 27.42 19.02 26.56 24.47 4.92 70.00 86.27 42.57 5.50 88.90 50.42 46.72 40.19
SLM 20.31 26.64 45.72 35.25 24.31 12.20 27.47 21.57 4.87 68.50 71.95 31.19 5.00 22.56 43.38 37.08 31.12
H2O 25.76 31.10 49.03 40.76 26.52 17.07 24.81 23.64 4.81 55.00 86.35 42.48 5.50 88.15 49.93 46.57 38.59
Ours 25.53 32.21 48.97 42.26 27.50 19.36 26.60 23.97 4.97 71.00 86.25 42.94 4.50 87.90 53.12 47.21 40.27
@@
Table 1: Performance comparison of  PyramidKV (Ours) with SnapKV (SKV), H2O, StreamingLLM (SLM) and FullKV (FKV) on LongBench for LlaMa-3-8B-Instruct and Mistral-7B-Instruct.  PyramidKV generally outperforms other KV Cache compression methods across various KV Cache sizes and LLMs. The performance strengths of  PyramidKV are more evident in small KV Cache sizes (i.e. KV Size = 128). Bold text represents the best performance.
@@
The evaluation results from LongBench[2] are shown in Table 1 and Figure 4. In Figure 4, we report the average score across datasets for 64, 96, 128, 256 and 512 case sizes. In Table 1, we report the results for two different KV cache sizes with 128 and 2048. These two sizes represent two distinct operational scenarios—the memory-efficient scenario and the performance-preserving scenario, respectively for a trade-off between memory and model performance. In Appendix E, we report the results of KV cache sizes with 64, 96, 128, 256, 512, 1024 and 2048.
Overall, PyramidKV preserves the performance with only 12% of the KV cache and it consistently surpasses other method across a range of KV cache sizes and different backbone models, with its performance advantages becoming particularly pronounced in memory-constrained environments where only about 0.7% of the KV cache from the prompt is retained. Upon examining specific tasks, PyramidKV demonstrates a notably superior performance on the TREC task, a few-shot question answering challenge. This suggests that the model effectively aggregates information from the few-shot examples, highlighting the potential for further investigation into in-context learning tasks.
5.6 PyramidKV Preserves the Long-Context Understanding Ability
We conduct the "Fact Retrieval Across Context Lengths" (Needle In A Haystack) experiment [23, 11] to evaluate the in-context retrieval capabilities of LLMs when utilizing various KV cache compression methods. For this purpose, we employ LlaMa-3-8B-Instruct and Mistral-7B-Instruct [18] as our base, with context lengths extending up to 8k and 32k tokens. We compared several KV cache compression techniques (PyramidKV, SnapKV [22], and H2O [34]) at cache sizes of 64 and full cache. The results are displayed in Figure 5 and Figure 6 (Additional results with varying KV cache sizes with Mistral-7B-Instruct-32k-v0.2 [18] and LlaMa-3-8B-8k are available in Appendix F). The results demonstrate that with only 64 KV cache retained, PyramidKV effectively maintains the model’s ability to understand short contexts (up to a context length of 6.8k), and experiences only a modest degradation in performance on longer contexts. In contrast, other KV cache compression methods significantly impair the performance of LLMs.
@@ 5.6 PyramidKV Preserves the Long-Context Understanding Ability
Figure 6: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in Mistral-7B-Instruct [18] with 32k context size in 64 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the token length.  PyramidKV mitigates the negative impact of KV cache compression on the long-context understanding capability of LLMs.
@@
5.7  PyramidKV Significantly Reduces Memory with Limited Performance Drop
@@
In this section, we study how sensitive the methods are with different sizes of KV cache. We report the KV cache memory reduction in Table 2. We evaluate the memory consumption of LLaMa-3-8B-Instruct. Specifically, we evaluate the memory consumption of all methods with a fixed batch size of 1, a sequence length of 8192, and model weights in fp16 format. We observe that PyramidKV substantially reduces the KV cache memory across different numbers of cache sizes.
@@
Table 2: Memory reduction by  PyramidKV. We compared the memory consumption between Llama-3-8B-Instruct with full KV cache and Llama-3-8B-Instruct compressed by  PyramidKV.
@@ 6 Conclusion
In this study, we investigate the intrinsic attention patterns of Large Language Models (LLMs) when processing long context inputs. Our empirical analysis leads us to discover the existence of Pyramidal Information Funneling in the attention across layers: it distributes information broadly in lower layers, increasingly concentrates it within specific contexts, and ultimately focuses on critical tokens in higher layers with massive activation or attention sink. Motivated by this discovery, we design a novel KV cache compression approach PyramidKV that utilizes this information flow pattern. It varies KV cache sizes across layers, uniquely designed to complement the attention behavior observed across different layers. By leveraging the Pyramidal Information Funneling into KV cache compression design, our method excels in memory-constrained settings, preserves long-context understanding ability, and significantly reduces memory usage with minimal performance trade-offs compared to the baselines.
7 Future Work
@@
Appendix A Limitations
Our experiments were limited to two base models: LLAMA-3-8B-Instruct and Mistral-7B-Instruct. While these models demonstrated consistent trends, the robustness of our findings could be enhanced by testing a broader array of model families, should resources permit. Additionally, our research was conducted exclusively in English, with no investigations into how these findings might be transferred to other languages. Expanding the linguistic scope of our experiments could provide a more comprehensive understanding of the applicability of our results globally.
Appendix B Details of Evaluation

### v4

@@ 5.1.3 Baselines
  • StreamingLLM (SLM) (Xiao et al., 2023) is an efficient framework that enables LLMs to accept infinite input length.
  • Heavy Hitter Oracle (H2O) (Zhang et al., 2024) is a KV cache compression policy that dynamically retains a balance of recent and Heavy Hitter (H2) tokens.
@@
    (Li et al., 2024) automatically compresses KV caches by selecting clustered important tokens for each attention head.
@@
5.2 Main Results
Figure 3: The evaluation results from LongBench (Bai et al., 2023) across 64, 96, 128 and 256 cache sizes at LLaMa-3-8B-Instruct (Left), Mistral-7B-Instruct (Middle) and LLaMa-3-70B-Instruct (Right). The evaluation metrics are the average score of LongBench across datasets.  PyramidKV outperforms H2O (Zhang et al., 2024), SnapKV (Li et al., 2024) and StreamingLLM (Xiao et al., 2023), especially in small KV cache sizes.
@@ 5.2 Main Results
LlaMa-3-8B-Instruct, KV Size = 64
@@
FKV 26.90 33.07 49.20 43.02 27.33 18.78 32.91 24.21 26.99 71.00 86.23 42.65 2.75 86.98 56.96 54.52 42.71
Mistral-7B-Instruct, KV Size = 64
SKV 16.94 17.17 39.51 36.87 22.26 15.18 14.75 20.35 21.45 37.50 84.16 37.28 4.50 61.13 42.40 38.44 30.72
SLM 15.01 13.84 28.74 30.97 24.50 13.42 13.25 19.46 19.17 35.50 76.91 29.61 4.67 27.33 38.71 35.29 25.60
H2O 18.19 19.04 37.40 30.18 22.22 13.77 16.60 21.52 21.98 37.00 81.02 38.62 5.00 66.03 43.54 40.46 30.88
Ours 20.91 20.21 39.94 33.57 22.87 15.70 17.31 21.23 21.41 54.00 81.98 36.96 3.58 60.83 44.52 37.99 32.19
@@
SKV 25.89 32.93 48.56 42.96 27.42 19.02 26.56 24.47 26.69 70.00 86.27 42.57 5.50 88.90 50.42 46.72 41.56
SLM 20.31 26.64 45.72 35.25 24.31 12.20 27.47 21.57 24.51 68.50 71.95 31.19 5.00 22.56 43.38 37.08 32.35
H2O 25.76 31.10 49.03 40.76 26.52 17.07 24.81 23.64 26.60 55.00 86.35 42.48 5.50 88.15 49.93 46.57 39.95
Ours 25.53 32.21 48.97 42.26 27.50 19.36 26.60 23.97 26.73 71.00 86.25 42.94 4.50 87.90 53.12 47.21 41.63
LlaMa-3-70B-Instruct, KV Size = Full
FKV 27.75 46.48 49.45 52.04 54.90 30.42 32.37 22.27 27.58 73.50 92.46 45.73 12.50 72.50 40.96 63.91 46.55
LlaMa-3-70B-Instruct, KV Size = 64
SKV 23.92 31.09 36.54 46.66 50.40 25.30 18.05 21.11 19.79 41.50 91.06 40.26 12.00 72.50 43.33 57.62 39.45
SLM 22.07 23.53 27.31 43.21 51.66 23.85 16.62 19.74 15.20 39.50 76.89 33.06 12.00 72.50 40.23 50.20 35.47
H2O 25.45 34.64 33.23 48.25 50.30 24.88 20.03 21.50 21.39 42.00 90.36 41.58 12.00 71.50 43.83 58.16 39.94
Ours 25.47 36.71 42.29 47.08 46.21 28.30 20.60 21.62 21.62 64.50 89.61 41.28 12.50 72.50 45.34 56.50 42.01
LlaMa-3-70B-Instruct, KV Size = 2048
SKV 26.73 45.18 47.91 52.00 55.24 30.48 28.76 22.35 27.31 72.50 92.38 45.58 12.00 72.50 41.52 69.27 46.36
SLM 26.69 41.01 35.97 46.55 52.98 25.71 27.81 20.81 27.16 69.00 91.55 44.02 12.00 72.00 41.44 68.73 43.96
H2O 27.67 46.51 49.54 51.49 53.85 29.97 28.57 22.79 27.53 59.00 92.63 45.94 12.00 72.50 41.39 63.90 45.33
Ours 27.22 46.19 48.72 51.62 54.56 31.11 29.76 22.50 27.27 73.50 91.88 45.47 12.00 72.50 41.36 69.12 46.55
@@
Table 1: Performance comparison of  PyramidKV (Ours) with SnapKV (SKV), H2O, StreamingLLM (SLM) and FullKV (FKV) on LongBench for LlaMa-3-8B-Instruct, Mistral-7B-Instruct and LlaMa-3-70B-Instruct.  PyramidKV generally outperforms other KV Cache compression methods across various KV Cache sizes and LLMs. The performance strengths of  PyramidKV are more evident in small KV Cache sizes (i.e. KV Size = 64).
@@
The evaluation results from LongBench (Bai et al., 2023) are shown in Table 1 and Figure 3. In Figure 3, we report the average score across datasets for 64, 96, 128, and 256 case sizes. In Table 1, we report the results for two different KV cache sizes with 64 and 2048. These two sizes represent two distinct operational scenarios—the memory-efficient scenario and the performance-preserving scenario, respectively for a trade-off between memory and model performance. In Appendix N, we report results of KV cache sizes with 64, 96, 128 and 2048.
@@
Figure 4: Results of the Fact Retrieval Across Context Lengths (“Needle In A HayStack”) test in LlaMa-3-70B-Instruct with 8k context size in 128 KV cache size. The vertical axis of the table represents the depth percentage, and the horizontal axis represents the length.
Overall, PyramidKV preserves the performance with only 12% of the KV cache and it consistently surpasses other method across a range of KV cache sizes and different backbone models, with its performance advantages becoming particularly pronounced in memory-constrained environments where only about 0.8% of the KV cache from the prompt is retained. Upon examining specific tasks, PyramidKV demonstrates a notably superior performance on the TREC task, a few-shot question answering challenge. This suggests that the model effectively aggregates information from the few-shot examples, highlighting the potential for further investigation into in-context learning tasks.
Notably, we initially observe the pyramidal attention patterns from the visualization analysis on the multi-document QA task (Figure 2), but the pyramid heuristic has demonstrated its effectiveness on a range of other LongBench tasks (e.g., single-document QA, In-Context Learning), suggesting its promising generalizability beyond multi-document QA.
The performance advantage of PyramidKV increases as the KV cache memory decreases. By focusing on optimizing budget allocation across layers, PyramidKV accurately allocates resources in memory-constrained scenarios, ensuring that retained information is effectively preserved to maintain model performance. Moreover, as in long bench results shown in Table 1, even in the performance-preserving scenario (i.e., KV cache size = 2048), PyramidKV improves the performance over baseline methods and even outperforms FullKV.
Among the 16 datasets, the tasks where our proposed method performs slightly worse than the baseline are mostly saturated (e.g., HotpotQA, Musique, etc under the LlaMa-3-8B-Instruct setting with KV Size = 64, as shown in Table 1). In these cases, our method is only marginally inferior to the baseline and remains competitive. Conversely, on tasks with greater potential for improvement (e.g., Qasper, MF-en, TREC, TriviaQA, etc under the same setting), our method significantly outperforms the baseline. Consequently, the overall average performance of our method surpasses that of the baselines. Notably, these tasks include several In-Context Learning tasks (i.e., TREC), our method enjoys best performance gain at In-Context Learning tasks.
5.3 Discussion and Insights
5.3.1 PyramidKV Preserves the Long-Context Understanding Ability
We conduct the "Fact Retrieval Across Context Lengths" (Needle In A Haystack) experiment (Liu et al., 2023a; Fu et al., 2024), which is a dataset designed to test whether a model can find key information in long input sequences, to evaluate the in-context retrieval capabilities of LLMs when utilizing various KV cache compression methods. For this purpose, we employ LlaMa-3-70B-Instruct as our base, with context lengths extending up to 8k. We compared several KV cache compression techniques (PyramidKV, SnapKV (Li et al., 2024), and H2O (Zhang et al., 2024)) at cache sizes of 128 and full cache. The results, presented in Figure 4  [^1] Additional results with 64, 96 and 128 KV cache sizes with LlaMa-3-8B-Instruct at 8k context length, LlaMa-3-70B-Instruct at 8k context length, and Mistral-7B-Instruct (Jiang et al., 2023) at 32k context length are available in Appendix P. The results demonstrate that with only 128 KV cache retained, PyramidKV effectively maintains the model’s ability to understand short contexts, and shows only modest degradation for longer contexts. In contrast, other KV cache compression methods significantly hinder the performance of LLMs. Notably, for the larger model (LlaMa-3-70B-Instruct), PyramidKV achieves 100.0 Acc. performance, matching the results of FullKV, thereby demonstrating its ability to preserve long-context comprehension with a substantially reduced KV cache. We adopt the haystack setting of haystack formed from a long corpus for the Needle In A Haystack task as Wu et al. (2024).
@@ 5.3.1 PyramidKV Preserves the Long-Context Understanding Ability
5.3.2  PyramidKV Significantly Reduces Memory with Limited Performance Drop
@@
In this section, we study how sensitive the methods are with different sizes of KV cache. We report the KV cache memory reduction in Table 2. We evaluate the memory consumption of LLaMa-3-8B-Instruct. Specifically, we evaluate the memory consumption of all methods with a fixed batch size of 1, a sequence length of 8192, and model weights in fp16 format. We observe that PyramidKV substantially reduces the KV cache memory across different numbers of cache sizes. We also present that the allocation strategy and score-based selection add minimal complexity in the inference phase as Appendix L.
@@
Table 2: Memory reduction effect and benchmark result by using PyramidKV. We conducted a comparison of memory consumption between the Llama-3-8B-Instruct model utilizing the Full KV cache and the Llama-3-8B-Instruct model compressed with the PyramidKV.
@@ 6 Conclusion
In this study, we investigate Pyramidal Information Funneling, the intrinsic attention patterns of Large Language Models (LLMs) when processing long context inputs. Motivated by this discovery, we design a novel KV cache compression approach PyramidKV that utilizes this information flow pattern. Our method excels in memory-constrained settings, preserves long-context understanding ability, and significantly reduces memory usage.
Appendix A Limitations
Our experiments were limited to three base models: LLAMA-3-8B-Instruct, LLAMA-3-70B-Instruct and Mistral-7B-Instruct. While these models demonstrated consistent trends, the robustness of our findings could be enhanced by testing a broader array of model families, should resources permit. Additionally, our research was conducted exclusively in English, with no investigations into how these findings might be transferred to other languages. Expanding the linguistic scope of our experiments could provide a more comprehensive understanding of the applicability of our results globally. Based on our results at LongBench and Needle-in-a-HayStack experiment, PyramidKV generally works decently in most of the language tasks (i.e., Single-Document QA, Multi-Document QA, Summerization, Few-Shot In-Context Learning, etc.). Although we observe that PyramidKV performs better in some tasks (i.e., Few-Shot In-Context Learning) compared with some other tasks (i.e., Summerization), we have not observed cases that the decoding result collapses at some tasks. This remains a new topic for future work to explore.
Appendix B Future Work
@@
Figure 5: Attention patterns of retrieval-augmented generation across layers in Mistral-7B-Instruct model [Jiang et al., 2023]
Figure 6: Attention patterns of retrieval-augmented generation across layers in Mixtral-8x7B-Instruct Mixture-of-Experts model.
Appendix C Related Work
Interpretation of LLMs
Prior research has shown that attention matrices in LLMs are typically sparse [Chen et al., 2024a, Xiao et al., 2023, Zhang et al., 2024], focusing disproportionately on a few tokens. For instance, Xiao et al. [2023] identified an “attention sink” phenomenon, where maintaining the Key and Value (KV) states of the first few tokens can substantially restore the performance of windowed attention, despite these tokens not being semantically crucial. Similarly, Sun et al. [2024] identified a “massive activations” pattern, where a minority of activations show significantly larger values than others within LLMs. Interestingly, these values remain relatively constant across different inputs and act as critical bias terms in the model.
KV Cache Compression
There has been a growing interest in addressing LLMs’ memory constraints on processing long context inputs. FastGen [Ge et al., 2023] introduces an adaptive KV cache management strategy that optimizes memory use by tailoring retention tactics to the specific nature of attention heads. This method involves evicting long-range contexts from heads that prioritize local interactions, discarding non-special tokens from heads focused on special tokens, and maintaining a standard KV cache for heads that engage broadly across tokens. SnapKV [Li et al., 2024] improves efficiency by compressing KV caches via selecting/clustering significant KV positions based on their attention scores. Heavy Hitter Oracle (H2O) [Zhang et al., 2024] implements a dynamic eviction policy that effectively balances the retention of recent and historically significant tokens, optimizing memory usage while preserving essential information. StreamingLLM [Xiao et al., 2023] enables LLMs trained on finite attention windows to handle infinite sequence lengths without fine-tuning, thus expanding the models’ applicability to broader contexts. LM-Infinite [Han et al., 2023] allows LLMs pre-trained with 2K or 4K-long segments to generalize to up to 200M length inputs while retaining perplexity without parameter updates.
Appendix D Pyramidal Information Funneling
Figure 5 and Figure 6 shows the attention patterns of one QA example over six different layers (i.e., 0, 6, 12, 18, 24, and 30) for Mistral-7B-Instruct model and Mixtral-8x7B-Instruct Mixture-of-Experts model. Figure 5 and Figure 6 demonstrate that the Pyramidal Information Funneling phenomenon is also evident in both the Mistral model and Mixtral model . The results reveal that, akin to Llama-like models, Mistral exhibit a progressively narrowing attention focus across layers. This supports the universality of the Pyramidal Information Funneling phenomenon across diverse model families. We hope this addresses your concern and underscores the generalizability of our findings.
Our analysis uniquely examines attention metrics across all transformer layers, from 0 to 30, leading to the discovery of a key phenomenon we term Pyramidal Information Funneling.
Lee et al. [2024] conducted a limited investigation into attention patterns, focusing only on the lower layer (layer 0) and a single upper layer (layer 18). While  Lee et al. [2024] noted that attention becomes more skewed in upper layers, it did not provide a fine-grained observation of attention patterns across all layers. In contrast, our study reveals several novel findings:
  • Localized Attention: We observe that attention progressively narrows its focus, targeting specific components within the input sequence.
  • Massive Attention Mechanism: In the upper layers, attention heavily concentrates on a small set of critical tokens. Notably, these tokens are not limited to the leading positions, as observed in  Lee et al. [2024], but also appear at regular intervals across the sequence. The discrepancy arises from differences in input settings, with  Lee et al. [2024] identifying massive attention only at the initial tokens.
These insights motivated us to propose a token-selection method based on the highest attention scores in the upper layers, rather than solely relying on tokens from earlier positions.
To the best of our knowledge,  Chen et al. [2024b] has not analyzed attention patterns across transformer layers.
Therefore, although  Lee et al. [2024] and  Chen et al. [2024b] are considered contemporaneous with our work, making a comparison unnecessary, the perspective of our observation is considered novel compared with  Lee et al. [2024] and  Chen et al. [2024b]. Moreover, although  Lee et al. [2024] also observed attention patterns, the method we proposed based on our observations is significantly different from  Lee et al. [2024], further highlighting the novelty of our work.
Appendix E Details of Proposed Method
Based on the pyramidal information funneling observed across different layers, PyramidKV consists of two steps: (1) Dynamically allocating different KV cache sizes/budgets across different layers; and (2) Selecting important KV vectors in each attention head for caching as  Figure 7.
Our decision to use an arithmetic sequence is driven by three key factors:
  • Alignment with Pyramidal Information Funneling Pattern: Empirical observations reveal a pyramidal information funneling pattern, where lower layers exhibit dispersed attention while higher layers concentrate on fewer tokens. Inspired by this, we adopt the arithmetic sequence design to align with this natural progression.
  • Superior Empirical Performance: Through extensive experimentation across diverse datasets, we compared various methods, including the arithmetic sequence and adaptive approaches. Results consistently showed that the arithmetic sequence method outperformed others.
  • Computational Efficiency: The arithmetic sequence method introduces minimal computational overhead compared to adaptive approaches, which require dynamically computing cache budgets across layers.
To perform KV cache eviction, we use torch.gather. Below, we outline the memory allocation and release process of torch.gather:
  • Index Selection: Identify the positions of the elements to extract from the input tensor.
  • Memory Location Calculation: Compute the specific memory locations of the elements to be extracted using the strides of the input tensor across each dimension.
  • Output Tensor Creation: Allocate memory to create a new output tensor and copy the selected elements to their corresponding positions in the output tensor.
  • Memory Management: Since torch.gather is not an in-place operation, it creates a new tensor to store the results, while the memory of the original input tensor is released.
The speed-up offered by PyramidKV is complementary to that achieved through tensor parallelism and pipeline parallelism, as these approaches are not mutually exclusive. PyramidKV can be seamlessly integrated with both tensor parallelism and pipeline parallelism.
Appendix F Details of Evaluation